# HN-MVTS: HyperNetwork-based Multivariate Time Series Forecasting

### v1

@@ Introduction
| 48 | 0.12550.0003 | 0.11840.0002 | 0.13770.0003 | 0.12200.0001 | 0.12700.0002 | 0.11910.0002 | 0.11620.0002 | 0.11570.0003 | 0.10910.0002 | 0.10760.0001 | |
| ECL | 96 | 0.14090.0004 | 0.13470.0004 | 0.17480.0011 | 0.14210.0003 | 0.15160.0007 | 0.13830.0005 | 0.13280.0003 | 0.13180.0002 | 0.13090.0002 | 0.12970.0002 |
| 192 | 0.15770.0010 | 0.15330.0009 | 0.19720.0012 | 0.16660.0008 | 0.17830.0009 | 0.17080.0006 | 0.15230.0005 | 0.15320.0005 | 0.15750.0005 | 0.15360.0004 | |
| 336 | 0.17200.0010 | 0.16760.0008 | 0.19860.0014 | 0.17470.0010 | 0.18040.0015 | 0.17500.0011 | 0.16690.0012 | 0.16810.0012 | 0.16900.0007 | 0.16370.0006 | |
| 48 | 0.27050.0022 | 0.26100.0017 | 0.26690.0044 | 0.27490.0049 | 0.26150.0017 | 0.25970.0013 | 0.25280.0063 | 0.26030.0049 | 0.28710.0005 | 0.28630.0004 | |
| ETTm1 | 96 | 0.30190.0023 | 0.28990.0021 | 0.30020.0014 | 0.29690.0010 | 0.29290.0004 | 0.28990.0003 | 0.28990.0019 | 0.28980.0020 | 0.30340.0037 | 0.31710.0039 |
| 192 | 0.33640.0010 | 0.33060.0009 | 0.34050.0008 | 0.33610.0008 | 0.34550.0010 | 0.33970.0011 | 0.32950.0034 | 0.33130.0029 | 0.33860.0087 | 0.35790.0105 | |
| 336 | 0.37170.0014 | 0.37210.0015 | 0.38200.0009 | 0.37590.0008 | 0.38240.0010 | 0.37600.0008 | 0.36570.0013 | 0.36690.0015 | 0.37670.0085 | 0.39460.0114 | |
| 48 | 0.12790.0003 | 0.12460.0003 | 0.13730.0004 | 0.13260.0003 | 0.12510.0002 | 0.12300.0001 | 0.13020.0005 | 0.13100.0009 | 0.13630.0006 | 0.13760.0011 | |
| ETTm2 | 96 | 0.16410.0003 | 0.16260.0004 | 0.18050.0005 | 0.17490.0004 | 0.16410.0007 | 0.16060.0003 | 0.16960.0009 | 0.17030.0010 | 0.18230.0013 | 0.18320.0015 |
| 192 | 0.21830.0003 | 0.21650.0003 | 0.23750.0007 | 0.23080.0006 | 0.22850.0010 | 0.22970.0014 | 0.23390.0012 | 0.23570.0016 | 0.25380.0010 | 0.24480.0009 | |
| 336 | 0.27460.0003 | 0.27300.0003 | 0.29060.0006 | 0.28480.0005 | 0.28570.0071 | 0.30750.0113 | 0.28850.0049 | 0.29030.0061 | 0.30550.0062 | 0.30550.0060 | |
| 48 | 0.15130.0002 | 0.14900.0001 | 0.10770.0003 | 0.10730.0003 | 0.13130.0007 | 0.11780.0003 | 0.11430.0004 | 0.10710.0002 | 0.08980.0001 | 0.08750.0001 | |
| PEMS03 | 96 | 0.18990.0002 | 0.18710.0002 | 0.13690.0003 | 0.13760.0004 | 0.18280.0014 | 0.20630.0049 | 0.14610.0021 | 0.13640.0020 | 0.11180.0005 | 0.10870.0003 |
| 192 | 0.20960.0011 | 0.20650.0009 | 0.15370.0021 | 0.15540.0028 | 0.20040.0034 | 0.16290.0029 | 0.17730.0008 | 0.17190.0006 | 0.13070.0003 | 0.12230.0002 | |
| 336 | 0.22650.0008 | 0.22270.0005 | 0.16540.0006 | 0.16450.0006 | 0.25050.0143 | 0.21350.0106 | 0.18990.0017 | 0.18620.0013 | 0.15040.0010 | 0.14810.0008 | |
| 48 | 0.16390.0008 | 0.15720.0005 | 0.11850.0006 | 0.11010.0003 | 0.10480.0004 | 0.10560.0005 | 0.13060.0014 | 0.11300.0011 | 0.09600.0004 | 0.08780.0003 | |
| PEMS04 | 96 | 0.20160.0012 | 0.19430.0009 | 0.13120.0007 | 0.12700.0007 | 0.17440.0009 | 0.16820.0010 | 0.16270.0132 | 0.14610.0066 | 0.11360.0015 | 0.10300.0011 |
| 192 | 0.22080.0014 | 0.21250.0012 | 0.15660.0043 | 0.13950.0039 | 0.18740.0075 | 0.16620.0034 | 0.18730.0068 | 0.16380.0029 | 0.12670.0024 | 0.11890.0018 | |
| 336 | 0.24440.0016 | 0.23900.0012 | 0.17600.0024 | 0.15190.0019 | 0.19850.0037 | 0.17860.0022 | 0.20050.0039 | 0.17410.0018 | 0.15330.0013 | 0.13330.0016 | |
| 48 | 0.14790.0005 | 0.14110.0005 | 0.09590.0002 | 0.09040.0001 | 0.13280.0004 | 0.10700.0003 | 0.09920.0001 | 0.08880.0002 | 0.06810.0001 | 0.06370.0001 | |
| PEMS07 | 96 | 0.18710.0020 | 0.17660.0016 | 0.11170.0006 | 0.10620.0005 | 0.12530.0067 | 0.14740.0106 | 0.12400.0063 | 0.10920.0024 | 0.07900.0002 | 0.07210.0001 |
| 192 | 0.20570.0018 | 0.19940.0011 | 0.13080.0008 | 0.12120.0010 | 0.13990.0023 | 0.14890.0026 | 0.14780.0012 | 0.12860.0010 | 0.09080.0003 | 0.08180.0002 | |
| 336 | 0.22870.0020 | 0.22000.0013 | 0.15330.0009 | 0.13260.0011 | 0.14420.0034 | 0.15450.0078 | 0.16190.0016 | 0.14150.0013 | 0.10300.0004 | 0.09280.0003 | |
| 48 | 0.18890.0035 | 0.16510.0027 | 0.11200.0014 | 0.10270.0009 | 0.12680.00011 | 0.12500.0013 | 0.12210.0038 | 0.10520.0015 | 0.08700.0003 | 0.07990.0002 | |
| PEMS08 | 96 | 0.26180.0105 | 0.22570.0074 | 0.14820.0040 | 0.12680.0028 | 0.22840.0035 | 0.23560.0042 | 0.15620.0014 | 0.13480.0008 | 0.11130.0003 | 0.09570.0003 |
| 192 | 0.30670.0063 | 0.27800.0026 | 0.22170.0027 | 0.18840.0013 | 0.23270.0029 | 0.18940.0015 | 0.21280.0019 | 0.19640.0008 | 0.14590.0016 | 0.11870.0009 | |
| 336 | 0.32730.0024 | 0.32520.0026 | 0.24960.0012 | 0.24770.0008 | 0.26940.0007 | 0.26400.0007 | 0.23430.0013 | 0.22310.0010 | 0.17700.0008 | 0.17690.0009 | |
| 48 | 0.13690.0025 | 0.11150.0007 | 0.11680.0008 | 0.11620.0007 | 0.11510.0004 | 0.11270.0005 | 0.11360.0007 | 0.11430.0008 | 0.12350.0004 | 0.11950.0003 | |
| Weather | 96 | 0.17330.0032 | 0.14250.0014 | 0.14900.0008 | 0.14750.0009 | 0.14800.0005 | 0.14290.0004 | 0.14480.0014 | 0.14800.0023 | 0.15590.0006 | 0.15050.0003 |
| 192 | 0.21670.0063 | 0.18570.0012 | 0.20000.0011 | 0.19480.0009 | 0.19410.0007 | 0.19140.0006 | 0.18650.0024 | 0.19030.0027 | 0.19920.0031 | 0.19950.0029 | |
| 336 | 0.26410.0042 | 0.23960.0028 | 0.24910.0010 | 0.24180.0011 | 0.25000.0009 | 0.24760.0006 | 0.24270.0003 | 0.24190.0002 | 0.25070.0006 | 0.24830.0003 |
@@
The training was performed on a server with 2 Nvidia A100 GPUs, 32 Intel Xeon Gold 6326 CPUs (2.90GHz) and 512 Gb of RAM. PyTorch 2.x was used to implement base models and our HN-MVTS. For the evaluation, we followed traditional settings (Nie et al. 2023; Wu et al. 2023): we set the input length of the loopback window to , and evaluate the results for horizon lengths of . For a fair comparison, we apply reversible instance normalization (Kim et al. 2021) for all models. We minimize the MSE with the Adam optimizer with learning rate 0.0001 and batch size 64. These hyperparameters let us achieve approximately the same results as reported in original papers (Zeng et al. 2023; Nie et al. 2023). The dimensionality of embeddings for HN-MVTS is set identical to the number of components in the time series: .
@@
We report average MSE for four horizon lengths and 5 random seeds in Table 2. Here, the statistically significant (under Wilcoxon signed-rank with confidence 0.95 over 5 seeds) model among HN-MVTS and the baseline is shown in bold. The Mean Absolute Error (MAE) and detailed training curves with dependence of testing MSE on epoch are presented in Appendix. As one can notice, in most cases, our HN-MVTS improves forecasting accuracy across all datasets and architectures, often by a substantial margin.
@@
Speaking about application to various architectures, first, the efficient and straightforward linear baseline, namely, DLinear (Zeng et al. 2023), benefits significantly from our method. For example, on the Weather dataset, MSE improves from 0.1369 to 0.1115 at horizon and from 0.2641 to 0.2396 at . As a result, Dlinear with our HN-MVTS became the top-performing solution for this dataset, though the vanilla Dlinear was initially the worst model. Similar gains are observed on ECL, PEMS04, and PEMS08, i.e., improvements are dataset-dependent and often practically meaningful (particularly in high-dimensional or strongly correlated datasets). These results confirm that HN-MVTS can enhance lightweight models by incorporating channel relationships. Second, as an MLP-based model, TSMixer (Chen et al. 2023) benefits from the added expressivity and channel specificity provided by our hypernetwork. For example, on the ECL dataset, MAE drops from 0.2453 to 0.2350 at , and from 0.3040 to 0.2876 at . Notably, TSMixer + HN-MVTS shows significant gains on PEMS07 and PEMS08, which involve high-dimensional traffic sensor data with strong inter-channel dependencies. Third, while already competitive, ModernTCN (Luo and Wang 2024) exhibits modest yet significant improvements when paired with HN-MVTS. On ETTm2 and PEMS datasets, the model becomes more robust at longer, typically more challenging horizons. Finally, even advanced Transformer-based models benefit from our approach. On the PEMS07 dataset at , PatchTST’s MSE decreases from 0.1619 to 0.1415, and iTransformer from 0.1030 to 0.0928. This demonstrates that our approach complements sophisticated attention mechanisms by generating adaptive, channel-aware output parameters. The visualization of HN-MVTS embeddings (Fig. 3) shows that channel embeddings learned for various base models are very similar and reflect the specifics of the dataset rather than the architecture of MVTS model.
@@
Regarding the dataset-specific observations, the Weather dataset shows the most significant improvement. For DLinear, the MSE at drops from 0.1733 to 0.1425 (an 18% relative improvement). ECL also sees noticeable improvement across horizons and architectures. For instance, TSMixer+HN-MVTS improves MSE from 0.1972 to 0.1666 at , showing that even a compact MLP benefits from our parameter-sharing scheme. Traffic datasets (PEMS), which contain hundreds of correlated sensor channels, particularly benefit from channel-specific modeling. At , HN-MVTS improves PatchTST on PEMS08 from 0.2343 to 0.2231, showing how our approach enhances scalability in high-dimensional input settings.
@@
Thus, our experiments demonstrate that while the absolute gains vary across datasets and backbones, the improvements are consistent in direction and sometimes substantial (e.g., Weather, traffic datasets), reflecting HN-MVTS’s role as a general plug-and-play mechanism rather than a dataset-specific optimization. Our HN-MVTS improves all models across nearly every dataset and horizon length, confirming the value of channel-aware, dynamic parameter generation. Moreover, the integration of our approach is model-agnostic: gains are achieved across all model types without architectural overhauls, underscoring the plug-and-play nature of our framework. Furthermore, we observe a resilience to horizon length . Performance gains persist across short and long horizons. Particularly for longer-term forecasts (e.g., ), where models typically degrade, HN-MVTS mitigates accuracy loss.
@@
To assess the computational overhead introduced by HN-MVTS, we measured the average training time per epoch for each model on all datasets and forecast horizons. Results are presented in Table 3. Despite introducing hypernetwork and learnable embeddings, the additional training cost remains modest. Across all models and datasets, the training time increase ranges from approximately 5% to 25%, depending on the model complexity and input size. The most lightweight models, such as DLinear and TSMixer, experience the most minor increases (e.g., DLinear on ECL at : from 15.31s to 17.24s, or 1̃2.6%). The relative increase is moderate given their baseline costs, even for more complex Transformer-based models like PatchTST and iTransformer. Even on high-dimensional datasets (e.g., PEMS07 with 883 channels), the added parameter cost is modest, and training time increases are negligible, demonstrating scalability. Thus, the HN-MVTS adds minimal training overhead, making it a practical enhancement for research and production scenarios. This efficiency and significant performance gains further strengthen the case for integrating HN-MVTS into a wide range of forecasting pipelines.
@@ Conclusion
In this work, we introduced HN-MVTS, a novel general-purpose framework for MVTS forecasting that combines the strengths of CI and CD modeling (Fig. 1). Unlike traditional CI models, which ignore inter-channel relationships, or CD models, which can be overly complex, our method leverages the concept of hypernetworks to generate channel-specific parameters in a lightweight and scalable manner. We enable the model to condition its predictions on channel similarity, allowing shared statistical strength across similar components while still preserving the robustness and efficiency of CI models. HN-MVTS contributes a training-time, architecture-agnostic generative prior for per-channel final-layer weights. Importantly, our approach is modular and architecture-agnostic: it can be seamlessly applied to a wide range of existing neural forecasting models (Fig. 2), including both CI (Dlinear, PatchTST) and CD (TSMixer, iTransformer). Through extensive empirical evaluations across multiple datasets, we demonstrated that HN-MVTS, in most cases, improves the accuracy of state-of-the-art forecasting models (Table 2). Moreover, these gains come at minimal computational cost, with only a modest increase in training time and parameter count (Table 3). The inference time is not affected as the hypernetwork may be removed at deployment phase after training the model.
Despite its flexibility and strong empirical results, HN-MVTS has several limitations. First, we assume that the last layer of a base forecasting model linearly transforms an input hidden state into the final output, so multi-layer hyperparameterization (e.g., generating weights beyond the last layer) might offer deeper insights into representation sharing. Second, although the method is architecture-agnostic, integrating it into highly specialized, non-standard forecasting architectures may require additional tuning or design adaptation. We did not consider the application of hypernetworks to non-neural network models (gradient boosting, statistical models, etc.), which are still widely used.

### v2

@@ Introduction
| 48 | 0.1255 | 0.1184 | 0.1377 | 0.1220 | 0.1270 | 0.1191 | 0.1162 | 0.1157 | 0.1091 | 0.1076 | |
| ECL | 96 | 0.1409 | 0.1347 | 0.1748 | 0.1421 | 0.1516 | 0.1383 | 0.1328 | 0.1318 | 0.1309 | 0.1297 |
| 192 | 0.1577 | 0.1533 | 0.1972 | 0.1666 | 0.1783 | 0.1708 | 0.1523 | 0.1532 | 0.1575 | 0.1536 | |
| 336 | 0.1720 | 0.1676 | 0.1986 | 0.1747 | 0.1804 | 0.1750 | 0.1669 | 0.1681 | 0.1690 | 0.1637 | |
| 48 | 0.2705 | 0.2610 | 0.2669 | 0.2749 | 0.2615 | 0.2597 | 0.2528 | 0.2603 | 0.2871 | 0.2863 | |
| ETTm1 | 96 | 0.3019 | 0.2899 | 0.3002 | 0.2969 | 0.2929 | 0.2899 | 0.2899 | 0.2898 | 0.3034 | 0.3171 |
| 192 | 0.3364 | 0.3306 | 0.3405 | 0.3361 | 0.3455 | 0.3397 | 0.3295 | 0.3313 | 0.3386 | 0.3579 | |
| 336 | 0.3717 | 0.3721 | 0.3820 | 0.3759 | 0.3824 | 0.3760 | 0.3657 | 0.3669 | 0.3767 | 0.3946 | |
| 48 | 0.1279 | 0.1246 | 0.1373 | 0.1326 | 0.1251 | 0.1230 | 0.1302 | 0.1310 | 0.1363 | 0.1376 | |
| ETTm2 | 96 | 0.1641 | 0.1626 | 0.1805 | 0.1749 | 0.1641 | 0.1606 | 0.1696 | 0.1703 | 0.1823 | 0.1832 |
| 192 | 0.2183 | 0.2165 | 0.2375 | 0.2308 | 0.2285 | 0.2297 | 0.2339 | 0.2357 | 0.2538 | 0.2448 | |
| 336 | 0.2746 | 0.2730 | 0.2906 | 0.2848 | 0.2857 | 0.3075 | 0.2885 | 0.2903 | 0.3055 | 0.3055 | |
| 48 | 0.1513 | 0.1490 | 0.1077 | 0.1073 | 0.1313 | 0.1178 | 0.1143 | 0.1071 | 0.0898 | 0.0875 | |
| PEMS03 | 96 | 0.1899 | 0.1871 | 0.1369 | 0.1376 | 0.1828 | 0.2063 | 0.1461 | 0.1364 | 0.1118 | 0.1087 |
| 192 | 0.2096 | 0.2065 | 0.1537 | 0.1554 | 0.2004 | 0.1629 | 0.1773 | 0.1719 | 0.1307 | 0.1223 | |
| 336 | 0.2265 | 0.2227 | 0.1654 | 0.1645 | 0.2505 | 0.2135 | 0.1899 | 0.1862 | 0.1504 | 0.1481 | |
| 48 | 0.1639 | 0.1572 | 0.1185 | 0.1101 | 0.1048 | 0.1056 | 0.1306 | 0.1130 | 0.0960 | 0.0878 | |
| PEMS04 | 96 | 0.2016 | 0.1943 | 0.1312 | 0.1270 | 0.1744 | 0.1682 | 0.1627 | 0.1461 | 0.1136 | 0.1030 |
| 192 | 0.2208 | 0.2125 | 0.1566 | 0.1395 | 0.1874 | 0.1662 | 0.1873 | 0.1638 | 0.1267 | 0.1189 | |
| 336 | 0.2444 | 0.2390 | 0.1760 | 0.1519 | 0.1985 | 0.1786 | 0.2005 | 0.1741 | 0.1533 | 0.1333 | |
| 48 | 0.1479 | 0.1411 | 0.0959 | 0.0904 | 0.1328 | 0.1070 | 0.0992 | 0.0888 | 0.0681 | 0.0637 | |
| PEMS07 | 96 | 0.1871 | 0.1766 | 0.1117 | 0.1062 | 0.1253 | 0.1474 | 0.1240 | 0.1092 | 0.0790 | 0.0721 |
| 192 | 0.2057 | 0.1994 | 0.1308 | 0.1212 | 0.1399 | 0.1489 | 0.1478 | 0.1286 | 0.0908 | 0.0818 | |
| 336 | 0.2287 | 0.2200 | 0.1533 | 0.1326 | 0.1442 | 0.1545 | 0.1619 | 0.1415 | 0.1030 | 0.0928 | |
| 48 | 0.1889 | 0.1651 | 0.1120 | 0.1027 | 0.1268 | 0.1250 | 0.1221 | 0.1052 | 0.0870 | 0.0799 | |
| PEMS08 | 96 | 0.2618 | 0.2257 | 0.1482 | 0.1268 | 0.2284 | 0.2356 | 0.1562 | 0.1348 | 0.1113 | 0.0957 |
| 192 | 0.3067 | 0.2780 | 0.2217 | 0.1884 | 0.2327 | 0.1894 | 0.2128 | 0.1964 | 0.1459 | 0.1187 | |
| 336 | 0.3273 | 0.3252 | 0.2496 | 0.2477 | 0.2694 | 0.2640 | 0.2343 | 0.2231 | 0.1770 | 0.1769 | |
| 48 | 0.1369 | 0.1115 | 0.1168 | 0.1162 | 0.1151 | 0.1127 | 0.1136 | 0.1143 | 0.1235 | 0.1195 | |
| Weather | 96 | 0.1733 | 0.1425 | 0.1490 | 0.1475 | 0.1480 | 0.1429 | 0.1448 | 0.1480 | 0.1559 | 0.1505 |
| 192 | 0.2167 | 0.1857 | 0.2000 | 0.1948 | 0.1941 | 0.1914 | 0.1865 | 0.1903 | 0.1992 | 0.1995 | |
| 336 | 0.2641 | 0.2396 | 0.2491 | 0.2418 | 0.2500 | 0.2476 | 0.2427 | 0.2419 | 0.2507 | 0.2483 |
@@
The training was performed on a server with 2x80Gb Nvidia A100 GPUs, Intel Xeon Gold 6326 CPUs (2.90GHz), and 512 GB of RAM. PyTorch 2.x was used to implement base models and our HN-MVTS. For the evaluation, we followed traditional settings (Nie et al. 2023; Wu et al. 2023): we set the input length of the loopback window to , and evaluate the results for horizon lengths of . For a fair comparison, we apply reversible instance normalization (Kim et al. 2021) for all models. We minimize the MSE with the Adam optimizer with a learning rate of 0.0001 and a batch size of 64. These hyperparameters enable us to achieve approximately the same results as those reported in the original papers (Zeng et al. 2023; Nie et al. 2023). The dimensionality of embeddings for HN-MVTS is set to be less than or equal to the number of components in the time series.
@@
We report average MSE for four horizon lengths and 5 random seeds in Table 2. Here, the statistically significant (under the Wilcoxon signed-rank test with confidence 0.95 over 5 seeds) model among HN-MVTS and the baseline is shown in bold. The Mean Absolute Error (MAE) and detailed training curves, with the dependence of testing MSE on epoch, are presented in Appendix. As one can notice, in most cases, our HN-MVTS improves forecasting accuracy across all datasets and architectures, often by a substantial margin.
@@
Regarding applications to various architectures, the efficient and straightforward linear baseline, namely DLinear (Zeng et al. 2023), benefits significantly from our method. For example, on the Weather dataset, MSE improves from 0.1369 to 0.1115 at horizon and from 0.2641 to 0.2396 at . As a result, DLinear, combined with our HN-MVTS, became the top-performing solution for this dataset, although the vanilla DLinear was initially the worst model. Similar gains are observed on ECL, PEMS04, and PEMS08, i.e., improvements are dataset-dependent and often practically meaningful (particularly in high-dimensional or strongly correlated datasets). These results confirm that HN-MVTS can enhance lightweight models by incorporating channel relationships.
Second, as an MLP-based model, TSMixer (Chen et al. 2023) benefits from the added expressivity and channel specificity provided by our hypernetwork. For example, on the ECL dataset, MAE drops from 0.2453 to 0.2350 at , and from 0.3040 to 0.2876 at . Notably, TSMixer + HN-MVTS shows significant gains on PEMS07 and PEMS08, which involve high-dimensional traffic sensor data with strong inter-channel dependencies. Third, while already competitive, ModernTCN (Luo and Wang 2024) exhibits modest yet significant improvements when paired with HN-MVTS. On ETTm2 and PEMS datasets, it becomes more robust at longer, typically more challenging horizons.
Finally, even advanced Transformer-based models benefit from our approach. On the PEMS07 dataset at , PatchTST’s MSE decreases from 0.1619 to 0.1415, and iTransformer from 0.1030 to 0.0928. This demonstrates that our approach complements sophisticated attention mechanisms by generating adaptive, channel-aware output parameters. The visualization of HN-MVTS embeddings (Fig. 3) shows that channel embeddings learned for various base models are very similar and reflect the specifics of the dataset rather than the architecture of the MVTS model.
@@
Regarding the dataset-specific observations, the Weather dataset shows the most significant improvement. For DLinear, the MSE at drops from 0.1733 to 0.1425 (an 18% relative improvement). ECL also sees noticeable improvement across horizons and architectures. For instance, TSMixer+HN-MVTS improves MSE from 0.1972 to 0.1666 at , showing that even a compact MLP benefits from our parameter-sharing scheme. Traffic datasets contain hundreds of correlated sensor channels, making them particularly suitable for channel-specific modeling. At , HN-MVTS improves PatchTST on PEMS08 from 0.2343 to 0.2231, showing how our approach enhances scalability in high-dimensional input settings.
@@
Thus, our experiments demonstrate that while the absolute gains vary across datasets and backbones, the improvements are consistent in direction and sometimes substantial (e.g., Weather, traffic datasets), reflecting HN-MVTS’s role as a general plug-and-play mechanism rather than a dataset-specific optimization. Our HN-MVTS improves all models across nearly every dataset and horizon length, confirming the value of channel-aware, dynamic parameter generation. Moreover, our approach is model-agnostic, yielding gains across all model types without requiring architectural overhauls, thereby underscoring the plug-and-play nature of our framework. Furthermore, we observe a resilience to horizon length . Performance gains persist across short and long horizons. Particularly for longer-term forecasts (e.g., ), where models typically degrade, HN-MVTS mitigates accuracy loss.
@@
To assess the computational overhead introduced by HN-MVTS, we measured the average training time per epoch for each model on all datasets and forecast horizons. Results are presented in Table 3. Despite introducing hypernetwork and learnable embeddings, the additional training cost remains modest. Across all models and datasets, the training time increase ranges from approximately 5% to 25%, depending on the model complexity and input size. The most lightweight models, such as DLinear and TSMixer, experience the most minor increases (e.g., DLinear on ECL at : from 15.31s to 17.24s, or 1̃2.6%). The relative increase is moderate given their baseline costs, even for more complex Transformer-based models like PatchTST and iTransformer. Even on high-dimensional datasets (e.g., PEMS07 with 883 channels), the added parameter cost is modest, and the increase in training time is negligible, demonstrating scalability. Thus, the HN-MVTS adds minimal training overhead, making it a practical enhancement for research and production scenarios. These efficiency gains and performance improvements further strengthen the case for integrating HN-MVTS into a wide range of forecasting pipelines.
@@ Conclusion
In this work, we introduced HN-MVTS, a novel general-purpose framework for MVTS forecasting that combines the strengths of CI and CD modeling (Fig. 1). Unlike traditional CI models, which ignore inter-channel relationships, or CD models, which can be overly complex, our method leverages the concept of hypernetworks to generate channel-specific parameters in a lightweight and scalable manner. We enable the model to condition its predictions on channel similarity, allowing shared statistical strength across similar components while still preserving the robustness and efficiency of CI models. HN-MVTS contributes a training-time, architecture-agnostic generative prior for per-channel final-layer weights. Importantly, our approach is modular and architecture-agnostic, allowing it to be seamlessly applied to a wide range of existing neural forecasting models (Fig. 2), including both CI (DLinear, PatchTST) and CD (TSMixer, iTransformer). Through extensive empirical evaluations across multiple datasets, we demonstrated that HN-MVTS, in most cases, improves the accuracy of state-of-the-art forecasting models (Table 2). These gains come at minimal computational cost, with only a modest increase in training time and parameter count (Table 3). The inference time is not affected, as the hypernetwork can be removed during the deployment phase after the model has been trained.
Despite its flexibility and strong empirical results, HN-MVTS has several limitations. First, we assume that the last layer of a base forecasting model linearly transforms an input hidden state into the final output; therefore, multi-layer hyperparameterization (e.g., generating weights beyond the last layer) may offer deeper insights into representation sharing. Second, although the method is architecture-agnostic, integrating it into highly specialized, non-standard forecasting architectures may require additional tuning or design adaptation. We did not consider the application of hypernetworks to non-neural network models (gradient boosting, statistical models, etc.), which are still widely used.